# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ 3.2 Existing KSD Tests with Tilted Kernels are Qualitatively Robust
Since the lack of robustness of existing KSD tests with stationary kernels arises because the Stein kernel is unbounded, a natural approach is to choose another kernel $k$ which prevents this from happening. We now show that this can be achieved using tilted kernels, a common approach to induce robustness for parameter estimation (Barp et al.,, 2019; Matsubara et al.,, 2022). Specifically, we will show in Theorem 3 that, with a tilted kernel, a small proportion of contamination in the data has negligible impact on the outcome of the standard KSD test.
We first show that tilted kernels give rise to bounded Stein kernels. The proof is deferred to Section A.2.
@@
Notably, the function $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$ is non-negative since $u_{p}$ is a reproducing kernel. Moreover, Barp et al., (2024, Theorem 9) shows that KSD with such tilted kernels still satisfies $P$-separation, namely $D(Q,P)=0\iff Q=P$, provided that $\|\mathbf{s}_{p}\|_{2}$ grows at most root-exponentially and the spectral density of the translation-invariant kernel $k$—which exists by Bochner’s Theorem (Berlinet and Thomas-Agnan,, 2004, Theorem 20)—is bounded away from zero on compact sets.
Most stationary kernels used in KSD tests (such as the IMQ or squared-exponential kernels) satisfy the conditions in Lemma 2. In this construction, the weighting function $w$ is used to counteract the growth of the score function $\mathbf{s}_{p}$. When the score grows polynomially like $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=\mathcal{O}(\|\mathbf{x}\|_{2}^{r})$ for some $r>0$, it suffices to choose $w(\mathbf{x})=(1+a^{2}\|\mathbf{x}\|_{2}^{2})^{-b}$ for any $a>0$ and $b\geq 1/(2r)$. This weight function is common in the frequentist parameter estimation literature (Barp et al.,, 2019) and the generalized Bayesian inference literature (Matsubara et al.,, 2022; Altamirano et al.,, 2023, 2024; Duran-Martin et al.,, 2024).
The following result states that, with a tilted kernel, the outcome of the standard KSD test will not be significantly affected by small proportions of contamination in the data. The proof is provided in Section A.3.
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma 2. Then, for any test level $\alpha\in(0,1)$ and any sequence $\epsilon_{n}=o(n^{-1/2})$, the following holds as $n\to\infty$,
[…]
@@
The LHS of the above convergence is the worst-case difference between the rejection probabilities of the standard KSD test with and without data corruptions. This result thus immediately implies that the standard KSD test is qualitative robust to any sequence of Huber’s models so long as the contamination ratio decays sufficiently fast as $\epsilon_{n}=o(n^{-1/2})$. The condition on $\epsilon_{n}$ allows the expected number of contaminated data $n\epsilon_{n}$ to grow with $n$ but requires the growth rate to be slow. In particular, it is met when the expected number of contaminated data is bounded, i.e., $\epsilon_{n}=\mathcal{O}(n^{-1})$. This condition is not an artifact of the proof: in Section B.1, we show empirically that this rate is tight, i.e., the tilted-KSD test is no longer qualitatively robust when $\epsilon_{n}=n^{-r}$ for any $r\leq 1/2$.
@@
The intuition behind Theorem 3 is that, when the conditions in Lemma 2 are met so that the Stein kernel $u_{p}$ is bounded, the impact of any single outlier on the test statistic $D^{2}(\mathbb{X}_{n})$ can be bounded. In contrast, under the setting of Theorem 1, where $k$ is translation-invariant and thus $u_{p}$ is unbounded, even a single outlier can drive $D^{2}(\mathbb{X}_{n})$ to infinity. This is the key motivation for using a tilted kernel to guarantee robustness: by choosing a suitable weighting function $w$, the Stein kernel can be made bounded, thereby mitigating the effect of outliers.
@@
The boundedness condition on $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}$ required of Lemma 2 and Theorem 3 bears similarity with the ones imposed for KSD-based robust estimation methods (Barp et al.,, 2024; Matsubara et al.,, 2022). For example, Barp et al., (2019, Proposition 7) showed that their KSD-based estimator is globally bias-robust assuming $\mathbf{x}\mapsto\|\mathbf{s}_{p}(\mathbf{x})\|_{2}\int_{\mathbb{R}^{d}}\|k(\mathbf{x},\mathbf{x}^{\prime})\mathbf{s}_{p}(\mathbf{x}^{\prime})\|_{2}Q(\mathop{}\!\mathrm{d}\mathbf{x}^{\prime})$ is bounded. This is slightly weaker than those we assume in Lemma 2 and Theorem 3, but also harder to verify in practice.
@@
Theorem 3 implies that the minimax separation boundary cannot decay faster than $n^{-1/2}$ when considering alternatives that include Huber-contamination models (see also Remark 1.3). This can be compared to existing results on the separation boundaries of tests based on Maximum Mean Discrepancy (MMD, Müller,, 1997; Gretton et al.,, 2012). MMD is a broader family of discrepancies that recovers KSD when the kernel is chosen to be the Stein kernel $u_{p}$ (Barp et al.,, 2024, Theorem 1). The separation rates of MMD tests have been studied extensively (Balasubramanian et al.,, 2021; Hagrass et al., 2024b, ), but these works assume the alternative $Q$ has a density with respect to either the Lebesgue measure or $P$, which does not hold in our contamination setting. An exception is Hagrass et al., 2024a , which does not make this assumption. However, their results are limited to the two-sample setting, where the target distribution $P$ is unknown and approximated by finite samples, again differing from our setup.
Although tilted kernels enforce qualitative robustness, they are not enough to guarantee quantitative robustness. Indeed, so long as $k$ is characteristic, the consistency of the standard KSD test implies that $\Pr_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}))\to 1$ whenever $Q\neq P$ Liu et al., 2016, Proposition 4.2; Schrab et al., 2022, Theorem 3.3, so KSD tests with either stationary or tilted kernels are not quantitatively robust to any neighborhoods strictly larger than the singleton set $\{P\}$. We provide numerical evidence in Section 5.1.
@@ 4 Robust Kernel Goodness-of-fit Tests for KSD-Ball Uncertainty Sets
1:Input: Data $\mathbb{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}$; model $P$; uncertainty radius $\theta$; bootstrap sample size $B$; test level $\alpha$.
@@
6:Compute the (non-squared) bootstrapped quantile $q_{B,1-\alpha}(\mathbb{X}_{n})$ by (5) and reject $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})>q_{B,1-\alpha}(\mathbb{X}_{n})$.
@@
where $\theta\geq 0$ and for $P\in\mathcal{P}(\mathbb{R}^{d})$ with density $p\in\mathcal{C}^{1}$, we define $\mathcal{B}^{\mathrm{KSD}}(P;\theta)\coloneqq\{Q:\;D(Q,P)\leq\theta\}$. The advantage of using a KSD-ball as the uncertainty set is that it lends naturally to a simple, tractable test that is robust to the contamination models we reviewed in Section 2.3. We first describe our novel test, which we call robust-KSD test, and show that it is quantitatively robust to KSD balls in the sense of Definition 2. We then discuss in Section 4.2 how to choose the uncertainty radius $\theta$ to incorporate other contamination models, such as Huber’s contamination models and density-band models.
@@ 4.1 A Robust KSD Test
where $D(\mathbb{X}_{n})$ is the square-root of the V-statistic (2). The test rejects $H^{\mathrm{C}}_{0}$ for large values of $\Delta_{\theta}(\mathbb{X}_{n})$. The decision threshold should be chosen to control the Type-I error rate for all possible $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, i.e., we require $\gamma=\gamma(\mathbb{X}_{n})$ so that $\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\Delta_{n}(\mathbb{X}_{n})>\gamma)\leq\alpha$. One approach to construct $\gamma$ is to use deviation inequalities for bounded functions as done in Gretton et al., (2012, Corollary 11). However, such approach tends to be overly conservative (Gretton et al.,, 2012). Instead, we propose a bootstrap procedure to construct a decision threshold. We will show that choosing $\gamma=q_{\infty,1-\alpha}(\mathbb{X}_{n})$, the square-root of the (population) bootstrap quantile defined in (4), gives the desired Type-I error control asymptotically. Our robust-KSD test therefore rejects $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})$. This is summarized in Algorithm 1.
The validity and consistency of our robust KSD test is stated in the following result, proved in Section A.5. In particular, it implies that our robust-KSD test is quantitatively robust to KSD balls in the infinite-sample limit.
@@
    $\limsup_{n\to\infty}\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)\cap\mathcal{P}(\mathbb{R}^{d};w)}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\;\leq\;\alpha\;.$
@@
    (Consistency) For any $Q\in\mathcal{P}(\mathbb{R}^{d};w)\backslash\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, it holds that
@@
The condition $w\in\mathcal{C}_{b}^{1}$ requires the weighting function $w$ and its gradient to be bounded. In particular, this holds for $w(x)=1$, in which case $k(\mathbf{x},\mathbf{x}^{\prime})=h(\mathbf{x}-\mathbf{x}^{\prime})$ reduces to an untilted stationary kernel such as IMQ or squared-exponential kernels. Theorem 4 also assumes the data-generating distribution $Q$ finitely integrates the fourth moment of the weighted score, i.e., $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}^{4}$. This automatically holds for any $Q\in\mathcal{P}(\mathbb{R}^{d})$ when $k$ is a tilted kernel satisfying the conditions in Lemma 2, in which case $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}$ becomes bounded.
@@
In practice, the decision threshold $q_{\infty,1-\alpha}$ is intractable, and we again use the Monte Carlo estimate $q_{B,1-\alpha}$, defined as the squared root of (5), as an approximation. Compared with the standard KSD test targeting the point null $H_{0}:Q=P$, our robust-KSD test uses the same bootstrap procedure to compute the decision threshold, but has a slightly different test statistic $\Delta_{\theta}(\mathbb{X}_{n})$ instead of $D^{2}(\mathbb{X}_{n})$, given by (2), to account for the composite null. Moreover, as $\theta\to 0$, the test statistic $\Delta_{\theta}$ approaches $D(\mathbb{X}_{n})$, the test statistic of the standard test, and the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ falls to the singleton $\{P\}$, so the robust-KSD test reduces to the standard test. Our robust KSD test can hence be viewed as a generalization of the standard KSD test to the composite hypotheses in (8). In particular, the robust KSD test is also qualitatively robust whenever the conditions on $k$ and $s_{p}$ from Theorem 3 hold; see Proposition 14 in Section A.3 for a formal statement and proof.
@@
Another advantage of our robust KSD test being a direct generalization of the standard test is that no extra computation is required to guarantee robustness. For a given $\theta\geq 0$, our robust test only requires a minor transformation of the test statistic of the standard test. It also has the same computational cost as the standard test, namely $\mathcal{O}(n^{2}d)$. However, extra computation is potentially needed in determining $\theta$, which might require optimizing the Stein kernel. This will be discussed in detail in Section 4.2.

### v4

@@ 3.2 Existing KSD Tests with Tilted Kernels are Qualitatively Robust
Since the lack of robustness of existing KSD tests with stationary kernels arises because the Stein kernel is unbounded, a natural approach is to choose another kernel $k$ which prevents this from happening. We now show that this can be achieved using tilted kernels, a common approach to induce robustness for parameter estimation (Barp et al.,, 2019; Matsubara et al.,, 2022). Specifically, we will show in Theorem˜3 that, with a tilted kernel, a small proportion of contamination in the data has negligible impact on the outcome of the standard KSD test.
We first show that tilted kernels give rise to bounded Stein kernels. The proof is deferred to Section˜A.2.
@@
Notably, the function $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$ is non-negative since $u_{p}$ is a reproducing kernel. Moreover, Barp et al., (2024, Theorem 9) shows that KSD with such tilted kernels still satisfies $P$-separation, namely $D(Q,P)=0\iff Q=P$, provided that $\|\mathbf{s}_{p}\|_{2}$ grows at most root-exponentially and the spectral density of the translation-invariant kernel kk—which exists by Bochner’s Theorem (Berlinet and Thomas-Agnan,, 2004, Theorem 20)—is bounded away from zero on compact sets.
Most stationary kernels used in KSD tests (such as the IMQ or squared-exponential kernels) satisfy the conditions in Lemma˜2. In this construction, the weighting function $w$ is used to counteract the growth of the score function $\mathbf{s}_{p}$. When the score grows polynomially like $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=\mathcal{O}(\|\mathbf{x}\|_{2}^{r})$ for some $r>0$, it suffices to choose $w(\mathbf{x})=(1+a^{2}\|\mathbf{x}\|_{2}^{2})^{-b}$ for any $a>0$ and $b\geq 1/(2r)$. This weight function is common in the frequentist parameter estimation literature (Barp et al.,, 2019) and the generalized Bayesian inference literature (Matsubara et al.,, 2022; Altamirano et al.,, 2023, 2024; Duran-Martin et al.,, 2024).
The following result states that, with a tilted kernel, the outcome of the standard KSD test will not be significantly affected by small proportions of contamination in the data. The proof is provided in Section˜A.3.
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma˜2. Then, for any test level $\alpha\in(0,1)$ and any sequence $\epsilon_{n}=o(n^{-1/2})$, the following holds as $n\to\infty$,
$\sup_{Q\in\mathcal{P}(P;\epsilon_{n})}\big{|}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n}^{\ast})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}\;\to\;0\;,$
@@
The LHS of the above convergence is the worst-case difference between the rejection probabilities of the standard KSD test with and without data corruptions. This result thus immediately implies that the standard KSD test is qualitative robust to any sequence of Huber’s models so long as the contamination ratio decays sufficiently fast as $\epsilon_{n}=o(n^{-1/2})$. The condition on $\epsilon_{n}$ allows the expected number of contaminated data $n\epsilon_{n}$ to grow with $n$ but requires the growth rate to be slow. In particular, it is met when the expected number of contaminated data is bounded, i.e., $\epsilon_{n}=\mathcal{O}(n^{-1})$. This condition is not an artifact of the proof: in Section˜B.1, we show empirically that this rate is tight, i.e., the tilted-KSD test is no longer qualitatively robust when $\epsilon_{n}=n^{-r}$ for any $r\leq 1/2$.
@@
The intuition behind Theorem˜3 is that, when the conditions in Lemma˜2 are met so that the Stein kernel $u_{p}$ is bounded, the impact of any single outlier on the test statistic $D^{2}(\mathbb{X}_{n})$ can be bounded. In contrast, under the setting of Theorem˜1, where $k$ is translation-invariant and thus $u_{p}$ is unbounded, even a single outlier can drive $D^{2}(\mathbb{X}_{n})$ to infinity. This is the key motivation for using a tilted kernel to guarantee robustness: by choosing a suitable weighting function $w$, the Stein kernel can be made bounded, thereby mitigating the effect of outliers.
@@
The boundedness condition on $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}$ required of Lemma˜2 and Theorem˜3 bears similarity with the ones imposed for KSD-based robust estimation methods (Barp et al.,, 2024; Matsubara et al.,, 2022). For example, Barp et al., (2019, Proposition 7) showed that their KSD-based estimator is globally bias-robust assuming $\mathbf{x}\mapsto\|\mathbf{s}_{p}(\mathbf{x})\|_{2}\int_{\mathbb{R}^{d}}\|k(\mathbf{x},\mathbf{x}^{\prime})\mathbf{s}_{p}(\mathbf{x}^{\prime})\|_{2}Q(\mathop{}\!\mathrm{d}\mathbf{x}^{\prime})$ is bounded. This is slightly weaker than those we assume in Lemma˜2 and Theorem˜3, but also harder to verify in practice.
@@
Theorem˜3 implies that the separation boundary of the KSD tests cannot decay faster than $n^{-1/2}$ when considering alternatives that include Huber-contamination models (see also Remark˜1.3). This can be compared to existing results on the separation boundaries of tests based on Maximum Mean Discrepancy (MMD, Müller,, 1997; Gretton et al.,, 2012). MMD is a broader family of discrepancies that recovers KSD when the kernel is chosen to be the Stein kernel $u_{p}$ (Barp et al.,, 2024, Theorem 1). The separation rates of MMD tests have been studied extensively (Balasubramanian et al.,, 2021; Hagrass et al., 2024b, ), but these works assume the alternative $Q$ has a density with respect to either the Lebesgue measure or $P$, which does not hold in our contamination setting. An exception is Hagrass et al., 2024a , which does not make this assumption. However, their results are limited to the two-sample setting, where the target distribution $P$ is unknown and approximated by finite samples, again differing from our setup.
Although tilted kernels enforce qualitative robustness, they are not enough to guarantee quantitative robustness. Indeed, so long as $k$ is characteristic, the consistency of the standard KSD test implies that $\Pr_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}))\to 1$ whenever $Q\neq P$ Liu et al., 2016, Proposition 4.2; Schrab et al., 2022, Theorem 3.3, so KSD tests with either stationary or tilted kernels are not quantitatively robust to any neighborhoods strictly larger than the singleton set $\{P\}$. We provide numerical evidence in Section˜5.1.
@@ 4 Robust Kernel Goodness-of-fit Tests for KSD-Ball Uncertainty Sets
1:Input: Data $\mathbb{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}$; target distribution $P$; uncertainty radius $\theta$; bootstrap sample size $B$; test level $\alpha$.
@@
6:Compute the (non-squared) bootstrapped quantile $q_{B,1-\alpha}(\mathbb{X}_{n})$ by (5).
7:Reject $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})>q_{B,1-\alpha}(\mathbb{X}_{n})$.
@@
where $\theta\geq 0$ and for $P\in\mathcal{P}(\mathbb{R}^{d})$ with density $p\in\mathcal{C}^{1}$, we define $\mathcal{B}^{\mathrm{KSD}}(P;\theta)\coloneqq\{Q:\;D(Q,P)\leq\theta\}$. The advantage of using a KSD-ball as the uncertainty set is that it lends naturally to a simple, tractable test that is robust to the contamination models we reviewed in Section˜2.3. We first describe our novel test, which we call robust-KSD test, and show that it is quantitatively robust to KSD balls in the sense of Definition˜2. We then discuss in Section˜4.2 how to choose the uncertainty radius $\theta$ to incorporate other contamination models, such as Huber’s contamination models and density-band models.
@@ 4.1 A Robust KSD Test
where $D(\mathbb{X}_{n})$ is the square-root of the V-statistic (2). The test rejects $H^{\mathrm{C}}_{0}$ for large values of $\Delta_{\theta}(\mathbb{X}_{n})$. The decision threshold should be chosen to control the Type-I error rate for all possible $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, i.e., we require $\gamma=\gamma(\mathbb{X}_{n})$ so that $\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}(\Delta_{n}(\mathbb{X}_{n})>\gamma)\leq\alpha$. One approach to construct $\gamma$ is to use deviation inequalities for bounded functions as done in Gretton et al., (2012, Corollary 11). However, such approach tends to be overly conservative (Gretton et al.,, 2012). Instead, we propose a bootstrap procedure to construct a decision threshold. We will show that choosing $\gamma=q_{\infty,1-\alpha}(\mathbb{X}_{n})$, the square-root of the (population) bootstrap quantile defined in (4), gives the desired Type-I error control asymptotically. Our robust-KSD test therefore rejects $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})$. This is summarized in Algorithm˜1.
The validity and consistency of our robust KSD test is stated in the following result, proved in Section˜A.5.2. In particular, it implies that our robust-KSD test is quantitatively robust to KSD balls in the infinite-sample limit.
@@
    $\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)\cap\mathcal{P}(\mathbb{R}^{d};w)}\limsup_{n\to\infty}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\;\leq\;\alpha\;.$
@@
    $\lim_{n\to\infty}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\;=\;1\;.$
@@
The condition $w\in\mathcal{C}_{b}^{1}$ requires the weighting function $w$ and its gradient to be bounded. In particular, this holds for $w(x)=1$, in which case $k(\mathbf{x},\mathbf{x}^{\prime})=h(\mathbf{x}-\mathbf{x}^{\prime})$ reduces to an untilted stationary kernel such as IMQ or squared-exponential kernels. Theorem˜4 also assumes the data-generating distribution $Q$ finitely integrates the fourth moment of the weighted score, i.e., $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}^{4}$. This automatically holds for any $Q\in\mathcal{P}(\mathbb{R}^{d})$ when $k$ is a tilted kernel satisfying the conditions in Lemma˜2, in which case $\|w(\mathbf{x})\mathbf{s}_{p}(\mathbf{x})\|_{2}$ becomes bounded.
@@
In practice, the decision threshold $q_{\infty,1-\alpha}$ is intractable, and we again use the Monte Carlo estimate $q_{B,1-\alpha}$, defined as the squared root of (5), as an approximation. Compared with the standard KSD test targeting the point null $H_{0}:Q=P$, our robust-KSD test uses the same bootstrap procedure to compute the decision threshold, but has a slightly different test statistic $\Delta_{\theta}(\mathbb{X}_{n})$ instead of $D^{2}(\mathbb{X}_{n})$, given by (2), to account for the composite null. Moreover, as $\theta\to 0$, the test statistic $\Delta_{\theta}$ approaches $D(\mathbb{X}_{n})$, the test statistic of the standard test, and the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ falls to the singleton $\{P\}$, so the robust-KSD test reduces to the standard test. Our robust KSD test can hence be viewed as a generalization of the standard KSD test to the composite hypotheses in (8). In particular, the robust KSD test is also qualitatively robust whenever the conditions on $k$ and $s_{p}$ from Theorem˜3 hold; see Proposition˜14 in Section˜A.3 for a formal statement and proof.
@@
Another advantage of our robust KSD test being a direct generalization of the standard test is that no extra computation is required to guarantee robustness. For a given $\theta>0$, our robust test only requires a minor transformation of the test statistic of the standard KSD test. It hence has the same computational cost as the standard test, namely $\mathcal{O}(n^{2}d)$. However, extra computation is potentially needed in determining $\theta$, as it might require optimizing the Stein kernel. This will be discussed in detail in Section˜4.2.
Remark 4.1 (Pointwise and uniform controls).
The Type-I error control shown in Theorem 4 is pointwise over the null distributions $Q$. For robust tests, a more desirable control is often a uniform one, whereby the supremum over $Q$ is taken before the limit over $n$ (Lehmann and Romano,, 2022, Chapter 11). For the proposed test, a uniform control can be shown if the bootstrapped quantile is replaced by the ground-truth quantile, say $q_{1-\alpha}^{\ast}(\mathbb{X}_{n})$. That is (see also Remark˜20.1 in the appendix),
$\limsup_{n\to\infty}\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)\cap\mathcal{P}(\mathbb{R}^{d};w)}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})>q_{1-\alpha}^{\ast}(\mathbb{X}_{n})\big{)}\;\leq\;\alpha\;.$
However, extending this to the bootstrapped quantile requires uniformly bounding the bootstrap approximation errors, which is non-trivial. We leave this as an open question for future research. In the literature, most kernel-based robust tests that offer uniform controls are two-sample tests using either permutation (Schrab and Kim,, 2024) or deviation bounds (Sun and Zou,, 2023). Yet, these approaches are either not applicable to our one-sample setting or too conservative. This is why our test uses bootstrapping. Another kernel robust test that uses bootstrapping is Key et al., (2025), but their Type-I error control is also pointwise. In Appendix˜D, we propose an alternative robust test that leverages deviation bounds in a manner similar to Sun and Zou, (2023) to achieve uniform controls, but at the expense of a lower test power.